# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ 3.3. Performance of Localized RETE Nets
Performance of a RETE net $(N,p)$ with respect to both execution time and memory consumption is largely determined by the effective size of a consistent configuration $\mathcal{C}$ for $(N,p)$, which we define as $|\mathcal{C}|_{e}:=\sum_{n\in V^{N}}\sum_{m\in\mathcal{C}(n)}|m|$. In this context, we define the size of a match $m:Q\rightarrow H$ as $|m|:=|m^{V}|+|m^{E}|=|V^{Q}|+|E^{Q}|$. It can then be shown that localization incurs only a constant factor overhead on the effective size of $\mathcal{C}$ for any edge-dominated host graph: [^4] For non-edge-dominated host graphs, the number of matches for (marking-sensitive) vertex input nodes may exceed the number of matches for related edge-input nodes. If matches for marking-sensitive vertex input nodes make up the bulk of intermediate results, localization then introduces an overhead on effective configuration size that cannot be characterized by a constant factor. However, we expect this situation to be rare in practice, since it requires very sparse host graphs.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q=query(p)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. Assuming that storing a match $m$ requires an amount of memory in $O(|m|)$ and storing an element from $\overline{\mathbb{N}}$ requires an amount of memory in $O(1)$, storing $\mathcal{C}^{\Phi}$ requires memory in $O(|\mathcal{C}|_{e})$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net for query graph $Q$, $\mathcal{C}$ a consistent configuration for $(N,p)$, and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for $(N^{\Phi},p^{\Phi})=localize(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot(|Q_{a}|+|Q|))$ steps, with $|Q_{a}|$ the average size of matches in $\mathcal{C}$ and $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@ 4.1. Localized RETE for Extended Queries
In the most basic case, a set of matches for an extended graph query $(Q,\psi)$ may be considered correct under a given relevant subgraph if it contains no matches that violate $\psi$ and all matches for $Q$ that satisfy the condition $\psi$ and involve elements from the relevant subgraph:
@@
(Correctness of Extended Query Results under Subgraphs) We say that a set of matches $M$ for the extended graph query $(Q,\psi)$ into a host graph $H$ is correct under a subgraph $H_{p}\subseteq H$ iff $\forall m\in M:m\models\psi\wedge\forall m\in\mathcal{M}^{Q}_{H}:(m\models\psi\wedge m(Q)\cap H_{p}\neq\emptyset)\Rightarrow m\in M$.
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the query $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$ along $a$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q^{\prime},\psi^{\prime})}$.
  • For a nested graph condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the query $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q,\psi^{\prime})}$.
  • For a nested graph condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS_{1}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q,\psi_{1})}$, the RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS_{2}=RPS([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$ with left dependency $[\ltimes]^{\Phi}_{1}$ and right dependency $p^{\Phi}_{(Q,\psi_{2})}$.
@@
$\ltimes_{2}$$\Phi$$\ltimes_{1}$$\Phi$$N^{\Phi}_{(Q,\psi_{2})}$$\Phi$$RPS_{2}$$\Phi$$N^{\Phi}_{Q}$$\Phi$$N^{\Phi}_{(Q,\psi_{1})}$$\Phi$$RPS_{1}$$\Phi$
@@
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. The set of matches given by the stripped result set $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then correct under $H_{p}$.
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the result of $order^{\Psi}$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort(RPS_{l})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q^{\prime},\psi^{\prime})})\circ[\ltimes]^{\Phi}$.
  • For a nested graph condition of the form $\psi=\neg\psi^{\prime}$, the result of $order^{\Psi}$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort(RPS_{l})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi^{\prime})})\circ[\rhd]^{\Phi}$.
  • Finally, for a nested graph condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, the result of $order^{\Psi}$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort(RPS_{1})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{1})})\circ[\ltimes]^{\Phi}_{1}\circ toposort(RPS_{2})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{2})})\circ[\ltimes]^{\Phi}_{2}$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. Executing $(N^{\Phi},p^{\Phi})$ via $O=order^{\Psi}(N^{\Phi})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ a graph query. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ and $\mathcal{C}^{\Phi}=execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$. The set of matches from $Q$ into $H$ given by $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then correct under $H_{p}$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. Assuming that storing a match $m$ requires an amount of memory in $O(|m|)$ and storing an element from $\overline{\mathbb{N}}$ requires an amount of memory in $O(1)$, storing $\mathcal{C}^{\Phi}$ requires memory in $O(|\mathcal{C}|_{e})$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot|Q_{max}|)$ steps, with $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$ and $|Q_{max}|$ the size of the largest graph among $Q$ and graphs in $\psi$.

### v2

@@ 3.3. Performance of Localized RETE Nets
Performance of a RETE net $(N,p)$ with respect to both execution time and memory consumption is largely determined by the effective size of a consistent configuration $\mathcal{C}$ for $(N,p)$, which we define as follows:
{defi}
(Effective Size of RETE Net Configurations) The effective size of a configuration $\mathcal{C}$ for a RETE net $(N,p)$ is given by $|\mathcal{C}|_{e}:=\sum_{n\in V^{N}}\sum_{m\in\mathcal{C}(n)}|m|$, where we define the size of a match $m:Q\rightarrow H$ as $|m|:=|m^{V}|+|m^{E}|=|V^{Q}|+|E^{Q}|$.
It can then be shown that localization incurs only a constant factor overhead on the effective size of $\mathcal{C}$ for any edge-dominated host graph: [^4] For non-edge-dominated host graphs, the number of matches for (marking-sensitive) vertex input nodes may exceed the number of matches for related edge-input nodes. If matches for marking-sensitive vertex input nodes make up the bulk of intermediate results, localization then introduces an overhead on effective configuration size that cannot be characterized by a constant factor. However, we expect this situation to be rare in practice, since it requires very sparse host graphs.
@@
(RETE net localization via $localize$ introduces only a constant factor overhead on effective configuration size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q=query(p)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
(RETE net localization via $localize$ introduces only a constant factor overhead on memory consumption.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. Assuming that storing a match $m$ requires an amount of memory in $O(|m|)$ and storing an element from $\overline{\mathbb{N}}$ requires an amount of memory in $O(1)$, storing $\mathcal{C}^{\Phi}$ requires memory in $O(|\mathcal{C}|_{e})$.
@@
(Execution time overhead introduced by RETE net localization via $localize$ depends on query graph size compared to average match size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net for query graph $Q$, $\mathcal{C}$ a consistent configuration for $(N,p)$, and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for $(N^{\Phi},p^{\Phi})=localize(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot(|Q_{a}|+|Q|))$ steps, with $|Q_{a}|$ the average size of matches in $\mathcal{C}$ and $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@ 4.1. Localized RETE for Extended Queries
In the most basic case, a set of matches for an extended graph query $(Q,\psi)$ may be considered correct under a given relevant subgraph if it contains no matches that violate $\psi$ and it contains all matches for $Q$ that satisfy the condition $\psi$ and involve elements from the relevant subgraph:
@@
(Correctness of Extended Query Results under Subgraphs) We say that a set of matches $M$ for the extended graph query $(Q,\psi)$ into a host graph $H$ is correct under a subgraph $H_{p}\subseteq H$ iff $M\subseteq\{m\in\mathcal{M}^{Q}_{H}|m\models\psi\}$ and $\{m\in\mathcal{M}^{Q}_{H}|m\models\psi\wedge m(Q)\cap H_{p}\neq\emptyset\}\subseteq M$.
@@
{defi}
(Target Result Sets of Advanced Marking-Sensitive RETE Nodes) Let $H$ be a host graph with relevant subgraph $H_{p}\subseteq H$, $(N^{\Phi},p^{\Phi})$ a containing marking-sensitive RETE net, and $\mathcal{C}^{\Phi}$ a configuration for $(N^{\Phi},p^{\Phi})$.
@@
  • The target result set of a marking-sensitive semi-join node $[\ltimes]^{\Phi}$ along a non-empty partial graph morphism $a:Q_{l}\rightarrow Q_{r}$ from some subgraph $Q_{p}\subseteq Q_{l}$ into $Q_{r}$ with marking-sensitive dependencies $n^{\Phi}_{l}$ and $n^{\Phi}_{r}$ with $Q_{l}=query(n^{\Phi}_{l})$ and $Q_{r}=query(n^{\Phi}_{r})$ is given by $\mathcal{R}^{\Phi}([\ltimes]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m_{l},\phi_{l})|(m_{l},\phi_{l})\in\mathcal{C}^{\Phi}(n^{\Phi}_{l})\wedge\exists(m_{r},\phi_{r})\in\mathcal{C}^{\Phi}(n^{\Phi}_{r}):m_{l}|_{Q_{p}}=m_{r}\circ a\}$.
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the query $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$ along $a$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q^{\prime},\psi^{\prime})}$.
  • For a nested graph condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the query $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q,\psi^{\prime})}$.
  • For a nested graph condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS^{\infty}_{1}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q,\psi_{1})}$, the RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS^{\infty}_{2}=RPS^{\infty}([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$ with left dependency $[\ltimes]^{\Phi}_{1}$ and right dependency $p^{\Phi}_{(Q,\psi_{2})}$.
Note that we include semi-join and anti-join nodes when computing join tree height for determining the value of marking filter nodes in request projection structures. Moreover, by $RPS^{\infty}$, we denote a request projection structure where the marking assignment node assigns a marking of $\infty$. While the marking assignment nodes in these request projection structures are technically redundant when considering the constructions in this article and could be removed, we include the nodes to mirror the structure of the request projection structures from Section 3.
@@
$\ltimes_{2}$$\Phi$$\ltimes_{1}$$\Phi$$N^{\Phi}_{(Q,\psi_{2})}$$\Phi$$RPS^{\infty}_{2}$$\Phi$$N^{\Phi}_{Q}$$\Phi$$N^{\Phi}_{(Q,\psi_{1})}$$\Phi$$RPS^{\infty}_{1}$$\Phi$
@@
The localized RETE net for the extended graph query in Figure 3 is visualized in Figure 8 along with a consistent configuration for the host graph in Figure 9. For local navigation structures and request projection structures, the displayed current result set is the current result set of the topmost union node respectively the marking assignment node of the structure.
Analogously to Figure 3, the localized RETE net essentially extends the RETE net from Figure 5 by a marking sensitive semi-join and a local navigation structure for matching the pattern $Q^{\prime}$. In addition, it contains a request projection structure that controls the computation of matches for $Q^{\prime}$ based on matches for $Q$ that touch the relevant subgraph.
Via this new request projection structure, the match $m_{1}$ for $Q$ triggers the extraction of the match $m_{1.3}$ for $Q^{\prime}$ in the local navigation structure $LNS^{c\rightarrow i}$. This ensures that $m_{1}$ is included in the result set of the semi-join node at the top of the net, which thereby provides a locally correct query result.
[…]
Figure 8. Localized RETE net for the example query in Figure 3 with a consistent configuration for the host graph in Figure 9
r1:Projp1:Pkgc1:Classf1:Fieldi1:Intfpececefeiep2:Pkgc2:Classf2:Fieldpecefe m1 m1.1 m1.1.1 m1.2 m1.2.1 m2 m1.3
Figure 9. Example host graph with relevant subgraph marked in bold
@@
(Consistent configurations for RETE nets localized via $localize^{\Psi}$ yield correct query results under the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. The set of matches given by the stripped result set $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then correct under $H_{p}$.
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the result of $order^{\Psi}$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort(RPS^{\infty}_{l})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q^{\prime},\psi^{\prime})})\circ[\ltimes]^{\Phi}$.
  • For a nested graph condition of the form $\psi=\neg\psi^{\prime}$, the result of $order^{\Psi}$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort(RPS^{\infty}_{l})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi^{\prime})})\circ[\rhd]^{\Phi}$.
  • Finally, for a nested graph condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, the result of $order^{\Psi}$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort(RPS^{\infty}_{1})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{1})})\circ[\ltimes]^{\Phi}_{1}\circ toposort(RPS^{\infty}_{2})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{2})})\circ[\ltimes]^{\Phi}_{2}$.
@@
(Execution of localized RETE nets via $order^{\Psi}$ yields consistent configurations.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. Executing $(N^{\Phi},p^{\Phi})$ via $O=order^{\Psi}(N^{\Phi})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
(Execution of RETE nets localized via $localize^{\Psi}$ yields correct query results under the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ a graph query. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ and $\mathcal{C}^{\Phi}=execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$. The set of matches from $Q$ into $H$ given by $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then correct under $H_{p}$.
@@
(RETE net localization via $localize^{\Psi}$ introduces only a constant factor overhead on effective configuration size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
(RETE net localization via $localize^{\Psi}$ introduces only a constant factor overhead on memory consumption.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. Assuming that storing a match $m$ requires an amount of memory in $O(|m|)$ and storing an element from $\overline{\mathbb{N}}$ requires an amount of memory in $O(1)$, storing $\mathcal{C}^{\Phi}$ requires memory in $O(|\mathcal{C}|_{e})$.
@@
(Execution time overhead introduced by RETE net localization via $localize^{\Psi}$ depends on query graph size compared to average match size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot|Q_{max}|)$ steps, with $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$ and $|Q_{max}|$ the size of the largest graph among $Q$ and graphs in $\psi$.